In positive economic models of the labor market, it is assumed that workers maximize utility (happiness)....

  1. In positive economic models of the labor market, it is assumed that

    workers maximize utility (happiness).

    firms maximize profits.

    resources are scarce.

    all of these are true.

2 points   

QUESTION 13

  1. Because of the principle of __________, marginal product of labor decreases as labor increases.

    economies of scale.

    diseconomies of scale.

    diminishing returns.

    constant returns to scale.

2 points   

QUESTION 14

  1. The unemployment rate is computed as

    the number of employed divided labor force.

    the number of unemployed divided the number of employed.

    the number of unemployed divided labor force.

    the sum of the numbers of unemployed and employed divided labor force.

2 points   

QUESTION 15

  1. Suppose the labor demand equation and labor supply equation: LS = 100 + W and LD =400 -2*W where LS = quantity supplied of labor, LD = quantity demanded for labor and W = wage. The equilibrium employment is _________ workers.

    50

    100

    200

    300

2 points   

QUESTION 16

  1. Suppose a firm is operating in both a perfectly competitive product market and perfectly labor market. The firm’s short run production is Q = L2; where Q is output and L is labor, expressed in millions. Marginal product of labor (MPL) = 2L and wage is 10. The price of the product is $ 2. Based on information above, the marginal revenue (MR) is

    $0.2

    $10

    $5

    $2

2 points   

QUESTION 17

  1. Suppose the marginal product of labor (MPL) = 100 – L, price of labor (wage) = $ 4 and the product price (P) = $ 2. In the short run, the optimal level of labor is

    96

    98

    99

    0

2 points   

QUESTION 18

  1. The labor force consists of

    All individuals aged 16 or older who are employed or unemployed.

    All individuals aged 16 or older who are employed.

    All individuals aged 16 or older who are unemployed.

    the discouraged

2 points   

QUESTION 19

  1. Suppose labor and capital are substitute and MRTSL,K = 1.5

    1 unit of capital can replace of 1.5 units of labor.

    1 unit of capital can replace of 1 unit of labor.

    1.5 unit of capital can replace of 1.5 units of labor.

    1 unit of labor can replace of 1.5 units of capital.

2 points   

QUESTION 20

  1. Suppose the labor demand equation and labor supply equation: LS = 100 + W and LD =400 -2*W where LS = quantity supplied of labor, LD = quantity demanded for labor and W = wage. The market clearing wage is ________ dollars.

    50

    100

    200

    300

2 points   

QUESTION 21

2

    5

    10

    50

2 points   

QUESTION 22

  1. Consider the following production function: Q = KL where Q = output, L = labor and K = capital. The marginal product of labor is given by MPL = K while the marginal product of capital is given by MPK = L.   If L = 10 and K= 5 the marginal rate of technical substitution between labor and capital is

    2

    5

    10

    1/2

2 points   

QUESTION 23

  1. The marginal product for labor is given (MP) = 3 – 0.02*L; price of the product is $100 and wage = 200. Based on information above, the optimal level of employment is in the short run is

    200

    100

    50

    25

2 points   

QUESTION 24

  1. 1-      Consider the following production function, Q = L2 + 20K. Marginal product of labor (MPL) = 2L and wage is $5. Marginal product of capital (MPK) = 20 and price of capital is $10. Then the MRTS at (L = 10 and K = 5), is equal to

    4

    2

    1

    1/2

2 points   

QUESTION 25

  1. 1-      Consider the following production function, Q = L2 + 20K. Marginal product of labor (MPL) = 2L and wage is $5. Marginal product of capital (MPK) = 20 and price of capital is $10. Then the relative price ratio is

    4

    2

    1

    1/2
Answer #1

Question - 12 8 - correct anders dizel maximize propita Question 13: Correct answers diminishing returns. Question 148 n an cI. Question 17:- correct answer - 98 In perfect competition : optimal laber is calculated by : - MRPL = MRX MPL 4 = 21100-L)W + 20 = 400-100 3 w = 300 or 300 Question alo- correct answer > 50 Q = KL = 5 x 10 K= 5 =10 = 50 Question 228- correct answeQuestion 248- correct answer -> 2 MRTS = L K => 10 - 2 5 -